OpenStudy (anonymous):

4x+16=8y+16 18y-22=3x+2

rebeccaxhawaii (rebeccaxhawaii):

add like terms

OpenStudy (anonymous):

how do you isolate each variable

rebeccaxhawaii (rebeccaxhawaii):

4x+16=8y+16 4x-8y=16-16 4x-8y=0

rebeccaxhawaii (rebeccaxhawaii):

18y-22=3x+2 -22-2=3x-18y -24=3x-18y

rebeccaxhawaii (rebeccaxhawaii):